OpenStudy (anonymous):

two angles are complementary. the sum of the measure of the first angle and half the second angle is 85 degrees. find the measures of the angles

OpenStudy (anonymous):

Angle A=80 degrees. Angle B= 10 degrees. You can solve this by simultaneous equations. A+B=90 A+B/2=85.

OpenStudy (anonymous):

x and y y=90-x =>x+y=90 x+y/2=85=>2x+y=170 x=80 y=10

OpenStudy (anonymous):

Thank you.